# NeuroTree: Hierarchical Functional Brain Pathway Decoding for Mental Health Disorders

### v1

@@ 3 Preliminaries
In this section, we formulate the mental disorder classification problem as a graph classification task, using fMRI networks constructed for each patient. We introduce the temporal brain activity measurements by giving $N$ patients with fMRI data. Let $X^{(i)}(t)\in\mathbb{R}^{v}$ denote the blood oxygen level-dependent (BOLD) signals at time point $t$ for subject $i$, where $v$ represents the number of regions of interest (ROIs). The complete time series for each subject can be represented as $X^{(i)}\in\mathbb{R}^{v\times T}$, where $T$ is the length of the time series. The foundation of brain network analysis lies in representing these temporal patterns through graph structures. For each subject, we construct a functional connectivity (FC) graph $G(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ represents the set of vertices (i.e., brain regions) and $\mathcal{E}$ denotes the set of edges (i.e., functional connections). The functional connectivity between regions is typically computed using Pearson or partial correlation methods. Based on these temporal correlation patterns, we derive both a static adjacency matrix $A^{s}$ from the complete time sequence and time-varying dynamic FC matrices $A^{d}(t)$. The static matrices capture averaged temporal dependencies, while the dynamic matrices preserve the temporal evolution of neural interactions [6, 7].
@@ 5.1 Optimal Weighted Tree Path
Given the pruned graph $\mathcal{T}(G)$, we examine paths that integrate node importance and high-order FC across multiple tree depths. Let $P=\{v_{1},v_{2},\dots,v_{k}\}$ represent a path of length $k-1$ in $\mathcal{T}(G)$, where each node $v\in P$ is associated with a scalar score $\mathcal{S}(v;\boldsymbol{\Theta})\in\mathbb{R}$, computed from Eq. 19. The edge set of $P$ is denoted by $E(P)=\{e_{i,j}\mid v_{i},v_{j}\in P(v_{i},v_{j})\in\mathcal{T}(G)\}$, where each edge $e_{ij}\in E(P)$ represent the path of neighbor connectivity. Let a single learnable node weight be defined:
$\mathcal{F}_{i,j}=\frac{1}{T}\sum_{t=1}^{T}\sigma\left(\sum_{v_{j}\in P}A_{i,j}^{d}(t)w_{j}(t)\right),$ (21)
where $A_{i,j}^{d}(t)$ represents the adjacency relationship between nodes $v_{i}$ and $v_{j}$ at time $t$, and $w_{j}(t)$ is the learnable weight for node $v_{j}$. The composite path weight $\mathcal{W}(P)$ is expressed as:
$\mathcal{W}(P)=\alpha\kern-8.5359pt$ $\underbrace{\sum_{v\in P}\mathcal{S}(v;\boldsymbol{\Theta})}_{\textbf{Node Score Contribution}}\kern-8.5359pt+(1-\alpha)\kern-5.69046pt\underbrace{\sum_{s=1}^{S}\sum_{(v_{i},v_{j})\in E(P)}\mathcal{F}_{i,j}^{(s)}}_{\textbf{High-Order FC Contribution}}.$ (22)
Here, $\mathcal{F}_{i,j}^{(s)}$ represents the $s$-th order FC between nodes $v_{i}$ and $v_{j}$ along the path and balance parameter $\alpha\in[0,1]$. To solve Eq. 22, we identify the path that minimizes $\mathcal{W}(P)$ by adjusting the parameter $\alpha$ to balance the contributions of the scoring function $\mathcal{S}(v;\boldsymbol{\Theta})$ and the connectivity strengths along the path. Additional experimental results can be found in Fig. 5. According to theorem 15 in the Appendix, the optimal path can be determined when $\alpha^{*}$ exist in the interval $[\alpha_{L},\alpha_{U}]$ [39].
Figure 2: Aggregation for hierarchical neighborhood paths Fig. (a) illustrates the brain tree structure with connectivity strengths derived from the pruned graph, providing an overall view of the hierarchical organization of functional connectivity. In Fig.(b), the zero-order path is depicted; here, aggregation initiates at the highest-scoring node, which integrates its immediate neighborhood by combining edges $C_{aj}$ and $C_{ah}$. Fig. (c) shows the aggregation of higher-order paths, where isolated nodes are connected along the shortest weighted paths, thereby capturing more complex connectivity patterns beyond immediate neighbors.
@@ 5.2 High-Order Tree Path Aggregation
$\mathcal{F}_{i,j}^{(s)}=\mathcal{F}_{i,j}^{(0)}+\sum_{p\in\mathcal{P}_{s}(i,j)}P_{s}(p)$ (23)
where $\mathcal{F}_{i,j}^{(0)}=A^{d}_{i,j}$ represents the direct average FC of a 0-th order path, and $\mathcal{P}_{s}(i,j)$ denotes the set of all possible s-order paths between nodes $i$ and $j$ in $\mathcal{T}(G)$. For each $s$-order path $p=(i,k_{1},k_{2},\ldots,k_{s},j)$, where all intermediate nodes are distinct ($k_{m}\neq k_{n}\;\forall m\neq n$), its contribution is calculated as $P_{s}(p)=\sum_{v\in p}\mathcal{F}_{v}$. For a path of order $s$, its extended connectivity at order $s+1$-th through the neighboring node $k$ can be computed as $\mathcal{F}_{i,j}^{(s+1)}=\sum_{k\in\mathcal{N}(i,j)}\left(\mathcal{F}_{i,k}^{(s)}+\mathcal{F}_{kj}^{(s)}\right)$, where $\mathcal{N}(i,j)$ represents the set of common neighbors between nodes $i$ and $j$, and $\mathcal{F}^{(s)}_{i,k}$ denotes the $s$-th order path connectivity from node $i$ to node $k$. This formulation allows us to capture direct connections and higher-order structural relationships within the brain network, as illustrated in Fig. 2.
@@ 5.3 Fine-Grained Tree Trunk
8:         Find $\nu_{\mathrm{start}}=\arg\max_{\nu\in\mathcal{K}_{j}}\mathcal{S}(\nu)$.
@@
$\mathcal{W}(P)=\alpha\sum_{v\in P}\mathcal{S}(v;\boldsymbol{\Theta})+(1-\alpha)\sum_{s=1}^{S}\sum_{(v_{i},v_{j})\in E(P)}\mathcal{F}_{i,j}^{(s)},$
10:         Let $\nu_{\mathrm{end}}$ maximize $\mathcal{W}(P)$, and define $P^{*}=\mathcal{P}(\nu_{\mathrm{start}},\nu_{\mathrm{end}})$.

### v2

@@ 3 Preliminaries
In this section, we formulate the mental disorder classification problem as a graph classification task, using fMRI networks constructed for each patient. We introduce the temporal brain activity measurements by giving $N$ patients with fMRI data. Let $X_{i}(t)\in\mathbb{R}^{v}$ denote the blood oxygen level-dependent (BOLD) signals at time point $t$ for subject $i$, where $v$ represents the number of regions of interest (ROIs). The complete time series for each subject can be represented as $X_{i}\in\mathbb{R}^{v\times T}$, where $T$ is the length of the time series. The foundation of brain network analysis lies in representing these temporal patterns through graph structures. For each subject, we construct a functional connectivity (FC) graph $G(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ represents the set of vertices (i.e., brain regions) and $\mathcal{E}$ denotes the set of edges (i.e., functional connections). The functional connectivity between regions is typically computed using Pearson or partial correlation methods. Based on these temporal correlation patterns, we derive both a static adjacency matrix $A^{s}$ from the complete time sequence and time-varying dynamic FC matrices $A^{d}(t)$. The static matrices capture averaged temporal dependencies, while the dynamic matrices preserve the temporal evolution of neural interactions [6, 7].
@@ 5.1 Optimal Weighted Tree Path
Given the pruned graph $\mathcal{T}(G)$, we examine paths that integrate node importance and high-order FC across multiple tree depths. Let $P=\{v_{1},v_{2},\dots,v_{k}\}$ represent a path of length $k-1$ in $\mathcal{T}(G)$, where each node $v\in P$ is associated with a scalar score $\mathcal{S}(v;\boldsymbol{\Theta})\in\mathbb{R}$, computed from Eq. 19. The edge set of $P$ is denoted by $E(P)=\{e_{ij}\mid v_{i},v_{j}\in P(v_{i},v_{j})\in\mathcal{T}(G)\}$, where each edge $e_{ij}\in E(P)$ represent the path of neighbor connectivity. Let a single learnable node weight be defined:
$\mathcal{F}_{ij}=\frac{1}{T}\sum_{t=1}^{T}\sigma\left(\sum_{v_{j}\in P}A_{ij}^{d}(t)w_{j}(t)\right),$ (21)
where $A_{ij}^{d}(t)$ represents the adjacency relationship between nodes $v_{i}$ and $v_{j}$ at time $t$, and $w_{j}(t)$ is the learnable weight for node $v_{j}$. The composite path weight $\mathcal{W}(P)$ is expressed as:
$\mathcal{W}(P)=\alpha\kern-8.5359pt$ $\underbrace{\sum_{v\in P}\mathcal{S}(v;\boldsymbol{\Theta})}_{\textbf{Node Score Contribution}}\kern-8.5359pt+(1-\alpha)\kern-5.69046pt\underbrace{\sum_{s=1}^{S}\sum_{(v_{i},v_{j})\in E(P)}\mathcal{F}_{ij}^{(s)}}_{\textbf{High-Order FC Contribution}}.$ (22)
Here, $\mathcal{F}_{ij}^{(s)}$ represents the $s$-th order FC between nodes $v_{i}$ and $v_{j}$ along the path and balance parameter $\alpha\in[0,1]$. To solve Eq. 22, we identify the path that minimizes $\mathcal{W}(P)$ by adjusting the parameter $\alpha$ to balance the contributions of the scoring function $\mathcal{S}(v;\boldsymbol{\Theta})$ and the connectivity strengths along the path. Additional experimental results can be found in Fig. 5. According to theorem 15 in the Appendix, the optimal path can be determined when $\alpha^{*}$ exist in the interval $[\alpha_{L},\alpha_{U}]$ [40].
Figure 2: Aggregation for hierarchical neighborhood paths Fig. (a) illustrates the brain tree structure with connectivity strengths derived from the pruned graph, providing an overall view of the hierarchical organization of functional connectivity. In Fig.(b), the zero-order path is depicted; here, aggregation initiates at the highest-scoring node, which integrates its immediate neighborhood by combining edges $e_{aj}$ and $e_{ah}$. Fig. (c) shows the aggregation of higher-order paths, where isolated nodes are connected along the shortest weighted paths, thereby capturing more complex connectivity patterns beyond immediate neighbors.
@@ 5.2 High-Order Tree Path Aggregation
$\mathcal{F}_{ij}^{(s)}=\mathcal{F}_{ij}^{(0)}+\sum_{p\in\mathcal{P}_{s}(i,j)}P_{s}(p)$ (23)
where $\mathcal{F}_{ij}^{(0)}=A^{d}_{ij}$ represents the direct average FC of a 0-th order path, and $\mathcal{P}_{s}(i,j)$ denotes the set of all possible s-order paths between nodes $i$ and $j$ in $\mathcal{T}(G)$. For each $s$-order path $p=(i,k_{1},k_{2},\ldots,k_{s},j)$, where all intermediate nodes are distinct ($k_{m}\neq k_{n}\;\forall m\neq n$), its contribution is calculated as $P_{s}(p)=\sum_{v\in p}\mathcal{F}^{(s)}_{v}$. For a path of order $s$, its extended connectivity at order $s+1$-th through the neighboring node $k$ can be computed as $\mathcal{F}_{ij}^{(s+1)}=\sum_{k\in\mathcal{N}(i,j)}\left(\mathcal{F}_{ik}^{(s)}+\mathcal{F}_{kj}^{(s)}\right)$, where $\mathcal{N}(i,j)$ represents the set of common neighbors between nodes $i$ and $j$, and $\mathcal{F}^{(s)}_{ik}$ denotes the $s$-th order path connectivity from node $i$ to node $k$. This formulation allows us to capture direct connections and higher-order structural relationships within the brain network, as illustrated in Fig. 2.
@@ 5.3 Fine-Grained Tree Trunk
8:         Find $\nu_{\mathrm{start}}=\arg\max\limits_{\nu\in\mathcal{K}_{j}}\mathcal{S}(\nu)$.
@@
$\mathcal{W}(P)=\alpha\sum_{v\in P}\mathcal{S}(v;\boldsymbol{\Theta})+(1-\alpha)\sum_{s=1}^{S}\sum_{(v_{i},v_{j})\in E(P)}\mathcal{F}_{ij}^{(s)}$ (24)
10:         Let $\nu_{\mathrm{end}}$ maximize $\mathcal{W}(P)$, and define $P^{*}=\mathcal{P}(\nu_{\mathrm{start}},\nu_{\mathrm{end}})$;